OpenStudy (anonymous):

how to divide the rational expression c^2-4/c / (c+2)

OpenStudy (anonymous):

\[\frac{c^2-\frac{4}{c}}{c+2}\]?

OpenStudy (anonymous):

or \[\frac{c^2-4}{\frac{c}{c+2}}\]

OpenStudy (anonymous):

\[(c ^{2} -4)/4 \div(c+2)\]

OpenStudy (gw2011):

I assume the problem is: [(c^2-4)/c]/(c+2) [(c^2-4)/c] [1/(c+2)] [(c-2)(c+2)/c] [1/(c+2)] (c-2)/c This is the answer to the problem